# Coherent Superconductor-Semiconductor Epitaxy for Integrated Quantum Electronics

### v2

@@ Acknowledgments
Figure S1: Comparison of this study against literature reports for ion implantation superconducting TC as a function of Hall carriers. The colored star, square, and triangular points were prepared via ion implantation [15, 2, 3] while the black circles (this study) are MBE-grown samples
@@
Figure S2: RHEED line cuts for superconducting hyperdoped Ge thin film. A Ge substrate $\langle 110\rangle$-type direction, taken at room temperature just before deposition of Ga:Ge film. B Initial surface of Ga:Ge film. C RHEED pattern after the first 10nm of Ga:Ge deposited. D After the Si spacer layer and initial growth of the second Ga:Ge layer. E End of Ga:Ge growth, before Si cap is deposited. We note a superimposed amorphous ring over a faint $\langle 110\rangle$-type direction diffraction pattern. This is consistent with previous observations that Ga-adsorption at low temperatures creates a floating layer at the surface from which the incorporated dopants are taken [17].
@@
Figure S4: Secondary-Ion Mass Specroscopy data for two samples investigated in this study, A PJS146 ($T_{\textup{c}}\sim$ 2.5 K) and B PJS147 ($T_{\textup{c}}\sim$ 3.5 K))
Figure S6: Comparison of the simulated fine structure derived from the CIF files of Ga:Ge interstitial system (DFT calculated system in Figure 2A of main text) and amorphous Ga metal [38] Note that these Fourier transforms have been phase corrected, for simplicity. The arrows highlight the relatively large difference in experimental bond length of the nearest-neighbor scattering pair and those derived in the simulations.
Figure S7: A Total density of states (DOS) calculated for the pure Ge crystal and the hyperdoped Ga:Ge shown in Fig. 2A of the main text, with Ga at interstitial (int.) and substitutional (sub.) lattice positions. The energies are aligned to the valence band maximum (VBM), set as the zero of the energy. The Fermi level in each case is indicated by a dashed vertical line. B Total and projected DOS of the hyperdoped Ga:Ge crystal with substitutional fine-structure.
Figure S9: Fermi surface obtained after shifting the Fermi level by 40 meV below the value shown in Figure 2B of the main text. A Contribution of bands i, ii, and iii to the 3D Fermi surface, with the color corresponding to the Fermi velocity. B (011) plane crossing the 3D Fermi surface. C 2D cut on this plane, showing the contribution of each band. The predicted narrow- or flat-band conditions are shown to be sensitive to the position of the Fermi level, as the pockets of high-mass carriers from bands ii and iii at the corners of the Brillouin Zone (R point) are shown to enlarge with a relatively minor shift in the Fermi level.
Figure S10: A Raman Scattering penetration depth ($\mathrm{d_{\mathrm{RS}}}$) as a function of the laser excitation energy, determined by considering the absorption coefficients [39] for both the incoming ($\alpha_{\mathrm{out}}$) and the outgoing ($\alpha_{\mathrm{in}}$) light: $\mathrm{d_{\mathrm{RS}}}=1/(\alpha_{\mathrm{in}}+\alpha_{\mathrm{out}})\approx 1/\alpha_{in}$. B Comparison of Raman spectra recorded from Ga:Ge and Ge control MBE films at different excitation energies. At longer excitation wavelengths (785 nm), both the control and target films exhibit a comparable Raman response. Probing the surface-sensitive information with shorter wavelengths, the differences between the vibrational structure of the films become clearer. The appearance of disorder-activated LA modes below 200 cm-1 (from both Ge-type and Si-type sublattices) in the scattering volume yields comparable scattering intensities under these conditions for the control and doped sample. We suggest the origins of these bands are a combination of (local) lattice strain, phonon confinement and (partial) satisfaction of the resonance conditions of the Ge crystal $E_{1}+\Delta_{1}$ gap [40]. Together, these features break the normal Raman scattering selection rules of a normal (001)-oriented bulk Ge film, i.e., resembling more the Raman spectra recorded using 785 nm laser light.
Figure S13: Evolution of the integrated GIWAXS signal recorded from the Ga:Ge film as a function of the grazing incidence angle ($\alpha$i: values inset). Here * identifies the calculated critical angle for the doped Ga:Ge layer at an X-ray beam energy of 20 keV. As the angle of incidence nears the critical angle of the doped layer, the Bragg reflections emerging begin to intensify. A combination of lateral steps (200 $\mu$m) made between each incident angle frame, mechanical jolting, and imperfect beam divergence means the critical angle conditions (waveguide type effect) is relatively really sensitive and results in the discontinuity of some profile features between frames.
@@
Figure S14: Integrated GIWAXS profile (qxyz) and structural refinement (Le Bail method) of the pure Ge epitaxial layer. Unstrained Ge crystal belongs to the high-symmetry cubic space group $Fd\overline{3}m$ (space group No. 227), characterized by equivalent lattice parameters (a=b=c) and a structural refinement of our undoped Si-Ge film aligns well with the reported bulk, cubic value (a = 5.657906$\pm$0.0000009 Å, Vol. = 181.1203 Å3)[41]
@@
Table S2: EXAFS fitting parameters. Values are extracted from fits made to spectra recorded at 295 K using a Ga substitutional model. The lower ranked paths (in italics) are omitted from the actual fit model to avoid overfitting. The errors provided in the parentheses align with the last significant figures of the variable fit values. Further fitting details can be found in the Methods section. The path type can be tracked as: SS - single scattering pair; DS - double scattering; R - rattle.
Path type Degen. Rank R(Å $\sigma^{3}$2) $\Delta$E0 (eV) S02 R-factor

### v3

@@ Methods
Hall bar device fabrication and measurement. A $\times5$ chip is first cleaned in acetone for 2 min, followed by isopropyl alcohol for 1 min, and then blow dried with compressed air. The chip is then dehydrated on a hotplate in a fume hood at 110C for 1 minute and then immediately placed into a spin coater. AZ 5214 photoresist (PR) is spun onto the chip at 4000 rpm for 60 s. Edge beads are removed using acetone and then the chip is placed back onto the hotplate at 110C for 1 min for the PR post-bake. The photolithography was conducted with a Süss MJB3 mask aligner using hard contact. The PR was exposed for 9 s, developed for 30 s in AZ 1:1 developer, and then rinsed in DI water for 15 s. The Hall bar was etched in an Oxford PlasmaPro System 100 Cobra inductively-coupled plasma (ICP) dry etcher system using a chlorine gas chemistry of 5 sccm Cl2, 5 sccm BCl3, 5 sccm Ar for 30 s at an ICP power of 500 W and a 75 W bias. This etch results in a step height of 140 nm as measured in an atomic force microscope, confirming a deep etch into the Ge substrate. The resist is stripped in AZ 400T PR stripper on a hotplate at 80C for 20 minutes, cleaned in isopropyl alcohol, and then blow dried. The device is wire bonded to a custom daughterboard and then cooled down in an Oxford Triton dilution refrigerator with a base temperature of 15 mK and a 14 T single-axis magnet.
Four-point transport measurements. Electrical characterization is done in an Oxford TeslatronPT He4 cryostat using a HelioxVT He3 probe insert with a base temperature of $\sim$ 270 mK (HelioxVT) and magnetic field capabilities up to 12 T. Measurements are collected in a similar fashion to Ref. Strohbeen et al. (2023), using a standard Van der Pauw wiring configuration on square pieces cleaved from near the center of each wafer. On-chip contacts are made via annealed In-Sn eutectic alloy at each of the four corners which are then wired to a daughterboard using gold wire.
First-principles calculations. First-principles calculations were performed using DFT as implemented in VASP Kresse and Hafner (1993); Kresse and Furthmüller (1996), except for the electron-phonon coupling calculations where we employed Quantum ESPRESSO Giannozzi et al. (2017). In VASP, projector augmented wave (PAW) pseudopotentials including semicore 3d electrons were adopted, with a plane-wave cutoff of 550 eV. A $7\times 7\times 7$ $\Gamma$-centered k-point mesh was used to sample the Brillouin zone of the conventional 8-atom unit cell of Ge (cubic diamond structure, space group $Fd\overline{3}m$). Structural optimizations were performed using the PBEsol exchange-correlation functional Perdew et al. (2008), which yields lattice parameters in excellent agreement with the experiment (see Table S1), and a convergence threshold of 0.005 eV/Å on all forces.
For the Ga-doped Ge structure with substitutional fine-structure, one host atom in the 8-atom unit cell is replaced by a Ga atom, corresponding to a doping concentration of 12.5%. Conversely, interstitial doping is modeled by placing one Ga atom at interstitial doping sites, representing a doping concentration of 11.1%. To model the effect of substrate clamping in Ga-doped Ge, the in-plane lattice parameters were fixed to the pristine Ge values, and only the out-of-plane lattice parameter was allowed to relax. K-edge spectra were simulated using the supercell core-hole method Karsai et al. (2018) within a $2\times 2\times 2$ supercell containing 64 atoms (72 atoms in the case of interstitial doping). The formation energy of a neutral Ga dopant in the Ge crystal is calculated as $E_{\textup{Form}}=E_{\textup{d}}-E_{\textup{p}}-\sum_{i}n_{i}\mu_{i}$, where $E_{\textup{d}}$ is the total energy of the structure containing the defect, $E_{\textup{p}}$ is the total energy of the pristine structure, $n_{i}$ is the number of atoms of species $i$ being added ($n>0$) or removed ($n<0$), and $\mu_{i}$ is the chemical potential. For Ga, this is the energy per atom of a Ga crystal (base-centered orthorhombic structure).
To calculate the electronic band structure and density of states, we adopted the MBJLDA meta-GGA functional Tran and Blaha (2009) based on the modified Becke-Johnson exchange functional. This functional has been shown to yield excellent results for the band structure and band gap of many semiconductors, including Ge Rödl et al. (2019); Fadaly et al. (2020), which standard semilocal DFT predicts erroneously to be metallic. Spin-orbit coupling was included. A denser $20\times 20\times 20$ $k$-mesh was employed to compute the density of states. To accurately determine the Fermi-level shifts in Ga-doped Ge, a $31\times 31\times 31$ $k$-mesh was used. Vibrational properties (shown in Fig. S15) were computed using finite differences Togo and Tanaka (2015) within supercells containing 216 atoms ($3\times 3\times 3$ supercells of the 8-atom unit cell), employing $3\times 3\times 3$ $k$-point meshes and displacement amplitudes of 0.015 Å.
We used Quantum ESPRESSO Giannozzi et al. (2017) to perform the electron-phonon coupling calculations, using norm-conserving pseudopotentials from the PseudoDojo library v0.4.1 van Setten et al. (2018) with standard accuracy and the PBEsol exchange-correlation functional, and a plane-wave cutoff of 80 Ry (1088 eV). We adopted the same model of substitutionally Ga-doped Ge, ignoring for simplicity the effect of substrate clamping. Relaxing the unit cell after Ga substitution results in a small lattice expansion of 0.09%. Spin-orbit coupling was neglected. We verified that the resulting Fermi level shift into the valence bands (0.96 eV) is consistent with the one obtained with the MBJLDA functional within VASP when spin-orbit coupling is neglected (0.93 eV). We calculated the Eliashberg spectral function $\alpha^{2}F(\omega)$ defined as Giustino (2017):
$\alpha^{2}F(\omega)=\frac{1}{N_{\textup{F}}N_{\mathbf{k}}N_{\mathbf{q}}}\sum_{mn\nu,\mathbf{k}\mathbf{q}}|g_{mn\nu}(\mathbf{k},\mathbf{q})|^{2}\,\delta(\varepsilon_{n\mathbf{k}}-\varepsilon_{\textup{F}})\delta(\varepsilon_{m\mathbf{k}+\mathbf{q}}-\varepsilon_{\textup{F}})\delta(\hbar\omega-\hbar\omega_{\mathbf{q}\nu}),$ (1)
where $N_{\textup{F}}$ is the density of states at the Fermi level, $N_{\mathbf{k}}$ and $N_{\mathbf{q}}$ are the total number of $\mathbf{k}$ and $\mathbf{q}$ points used to sample the Brillouin zone, $\varepsilon_{n\mathbf{k}}$ is the electron energy with band index $n$, $\varepsilon_{\textup{F}}$ is the Fermi energy, and $\hbar\omega_{\mathbf{q}\nu}$ is a phonon energy with momentum $\mathbf{q}$ and branch $\nu$. $g_{mn\nu}(\mathbf{k},\mathbf{q})$ are the electron-phonon matrix elements Giustino (2017) calculated using density-functional perturbation theory. The total electron-phonon coupling strength $\lambda$ is then obtained as $\lambda=2\int d\omega\frac{\alpha^{2}F(\omega)}{\omega}$. Using the semi-empirical Allen-Dynes-McMillan equation, the superconducting critical temperature $T_{\textup{c}}$ is estimated as McMillan (1968); Allen and Dynes (1975):
$k_{\textup{B}}T_{\textup{c}}=\frac{\hbar\omega_{\textup{log}}}{1.2}\,\mbox{exp}\left[\frac{-1.04(1+\lambda)}{\lambda-\mu^{\ast}(1+0.62\lambda)}\right],$ (2)
where $k_{\textup{B}}$ is Boltzmann’s constant, $\omega_{\textup{log}}$ is the logarithmic average of the phonon frequencies, given by $\omega_{\textup{log}}=\mbox{exp}\left[\frac{2}{\lambda}\int_{0}^{\infty}d\omega\frac{\alpha^{2}F(\omega)}{\omega}\log\omega\right]\!,$ and $\mu^{\ast}$ is the Coulomb pseudopotential, here set to 0.1 Allen and Dynes (1975). To converge the Brillouin-zone integrals in Eq. (1), we employed a $20\times 20\times 20$ and $5\times 5\times 5$ $k$ and $q$-mesh, respectively, and a Gaussian broadening of 0.04 Ry for the electronic double-delta functions.
Scanning transmission electron microscopy. Cross-sectional specimens for scanning transmission electron microscopy (S/TEM) analysis were prepared using a focused ion beam (FIB) system (FEI Helios NanoLab 600 DualBeam). To minimize surface damage, the samples were thinned and polished with Ga ions at 5 kV. S/TEM imaging and energy-dispersive X-ray spectroscopy (EDS) were conducted using a Thermo Fisher Scientific Themis-Z microscope. The instrument operated at an accelerating voltage of 300 kV with a semi-convergence angle of 20 mrad. HAADF-STEM images were collected using a 64-200 mrad high-angle annular dark field (HAADF) detector. High-resolution images were acquired through multiple rapid scans (2048 $\times$ 2048 pixels, 200 ns per pixel), which were combined to improve the signal-to-noise ratio. EDS analysis was performed with the Super-X EDS detector, and the elemental mappings are presented as net count images.
Raman Spectroscopy. Raman spectra were recorded from the thin film surface in a backscattering configuration, using an integrated Edinburgh confocal micro-Raman instrument (RM5). Excitation was provided by one of three solid-state laser lines (532 nm, 638 nm and 785 nm) and focused on the surface using a long working lens (Olympus, 10x, 0.4 NA). Dispersion was achieved through a 1200 g/mm diffraction grating (spectral resolution 0.2 cm-1), instrument calibration was verified through checking the position of the Si band at $\pm$520.7 cm-1, and spectra were recorded using a thermoelectric CCD detector. Changes to the local vibrational structure of the hyperdoped film are first confirmed through the evaluation of the excitation energy dependence of the Raman back-scattering signals (Figure S12). At longer excitation wavelengths (785 nm), both the control and target films exhibit comparable Raman response, while more surface-sensitive information is recovered at shorter wavelengths (532 nm) to resolve differences in the thin films (Figure 3f).
Synchrotron-based GISAXS/GIWAXS. To identify the crystalline phases and microstructure of films, synchrotron-based grazing incidence wide-angle X-ray scattering (GIWAXS) data were collected at the small angle X-ray scattering/wide angle X-ray scattering (SAXS/WAXS) beamline at the Australian Synchrotron Kirby et al. (2013): 2D scattering patterns were recorded using a wavelength of either 1.340375 Å (9.25 keV) or 0.619924 Å (20 keV), using an energy-selective Pilatus 2M detector. The measurements made using a 20 keV beam energy were to increase the observed Q-range for the structural refinements, using an energy threshold of 15 keV to suppress the X-ray fluorescence signals arising from the excited K-edges of Ge (11.1 keV) and Ga (10.37 keV). Experiments at 9.25 keV had no such issues. Two different sample-to-detector distances were employed to record both SAXS (5.5 m) and WAXS (0.45 m) profiles, each calibrated using a silver behenate reference standard. The sample and detector were enclosed in a vacuum chamber to suppress air scatter. Scattering patterns were measured as a function of the angle of incidence, with data shown acquired with an angle of incidence near the critical angle (for a given X-ray energy) to maximize scattering intensity from the sample. All collected 2D images were azimuthally integrated using PyFAI Ashiotis et al. (2015) and processed using a custom Python routine. The resulting unit cell models were refined using the La Bail method implemented in Fullprof Rodríguez-Carvajal (1993), a comprehensive analysis software. In line with our other characterization techniques, our GIWAXS experiments failed to detect Ga droplet formation or segregation within the film. Due to the relatively small scattering volume of the Si layers, no additional peaks are indexed to this phase either.
Synchrotron-based EXAFS experiments. Ga K-edge X-ray absorption spectra (XAS) were captured at the Australian Synchrotron (ANSTO in Clayton, Victoria) MEX1 beamline in a similar fashion to previously reported experiments Zhang et al. (2024). A double-crystal Si (111) monochromator equipped with focusing optics was used to reduce harmonic content while producing excitation energy. An inline Ga2O3 reference (the initial peak of the first derivative occurs at 10.3687 keV) was used to calibrate the monochromator at Ga-K absorption edge with E0 set to its known value of 10.3671 eV. All acquisitions were performed on thin film materials in a fluorescent mode. Brief XAS scans of specific samples were conducted to verify the stability of the materials under at least several minutes of X-ray exposure before the full acquisitions. Samples were measured with varying energy intervals over the pre-edge (5 eV) and the XANES region (0.25), with 0.035 intervals in $k$-space over the EXAFS. At low-$k$ an integration time of 1 s was employed per interval, with longer integration times weighted toward high-$k$ portions of the spectrum, up to a maximum value of 9 in $k$-space (max of 2 s). Multiple scans were collected at different positions of the sample. Data processing including background subtraction, scan averaging, edge-height normalization, and rebinning was performed based on the Athena program Ravel and Newville (2005). For normalization of energy spectra and removal of background, the pre-edge range was set between -150 eV and -30 eV, while the normalization range spanned from 80 to 300 eV post-edge. The order of normalization was designated as 3. For the presented EXAFS data, the $k$-weight was configured as 2, and the $k$-range for the forward Fourier transform was defined from 0.8 to 8.5.
EXAFS fine-structure modeling. Fits were conducted in the ARTEMIS, part of the IFEFFIT software package Ravel and Newville (2005) using a similar procedure to previous studies Zhang et al. (2024). The scattering paths used to evaluate and model the data were derived from the optimized crystal structures with different defects. Using these input structures to derive the scattering paths, the only agreeable fine structure accounting for the dominant EXAFS signal arising from an R fitting window of 1.8 to 5.1 Å emerged using the calculated Ga substituted structure. To simulate the Ga substituted Ge crystal fine structure, the input paths out to 5 Å are determined using the DFT-derived Ga substitutionally doped Ge structure and the FEFF package Ravel and Newville (2005), with a distance fuzz = 0.030 Å. This generated 6 unique scattering pairs centered on the Ga substitutional site, with some other lower ranked paths omitted from the actual model to avoid over-fitting (see Table S2). A common scaling factor has been used on all path distance corrections to account for the difference between the DFT simulated unit cell volume and the experiment crystal volume at room temperature. A phase correction has been made to Figure 2d in the main text using the most intense Ga-Ge single scattering path, near 2.45 Å. The data were ultimately fit using 18 variables and 26 independent points, resulting in a reasonable R-factor of 0.015. The fitting was conducted in R-space using multiple k-weightings, with k = 1, 2 and 3. Errors of individual fit parameters were determined using ARTEMIS to take into account the correlations between parameters and known parameters without error estimates were fixed during the fit.
[…]
Supplemental Information
S1 State of hyperdoping in germanium
@@
Figure S1: a Comparison of this study against literature reports for superconducting $T_{\textup{c}}$ as a function of Hall carriers in Ga-doped germanium. The colored star, square, and triangular points were prepared via ion implantation Hermannsdörfer et al. (2009); Prucnal et al. (2019); Sardashti et al. (2021) while the black circles (this study) and black diamond Strohbeen et al. (2023) are MBE-grown samples. b Representative AFM image of the surface of the sample circled in red in a. We observe an RMS roughness of roughly 1.65 nm, a significant improvement over previous MBE reports Strohbeen et al. (2023). The inset presents the normalized height distribution, $\rho$, across the entire 15 $\mu m\times$15 $\mu m$ image.
S2 Material optimization
@@
Figure S2: RHEED line cuts for superconducting hyperdoped Ge thin film. a Ge substrate $\langle 110\rangle$-type direction, taken at room temperature just before deposition of Ga:Ge film. b Initial surface of Ga:Ge film. c RHEED pattern after the first 10nm of Ga:Ge deposited. d After the Si spacer layer and initial growth of the second Ga:Ge layer. e End of Ga:Ge growth, before Si cap is deposited. We note a superimposed amorphous ring over a faint $\langle 110\rangle$-type direction diffraction pattern. This is consistent with previous observations that Ga-adsorption at low temperatures creates a floating layer at the surface from which the incorporated dopants are taken Kesan et al. (1991).
Film growth is monitored using reflection high-energy electron diffraction (RHEED) along a $\langle 110\rangle$ cut. Representative RHEED images throughout the growth steps are presented in Figure S2. The haziness observed in Figure S2b-e is caused by a surface floating layer of Ga that forms during the growth process Kesan et al. (1991). The Ga shutter is closed during Si deposition, however, we note that the floating layer slightly incorporates into the Si layer as indicated by the increased intensity in the specular RHEED reflection in Figure S2d compared to Figure S2c. To reduce parasitic heating of the substrate, the hot Ga crucible was backed up within the vacuum chamber by 4 in to reduce radiative heating of the growth surface when the shutter is open.
Figure S3: Superconducting $T_{\textup{c}}$ comparing the effects of silicon as a function of hole concentration and growth temperature. The circular data points are the samples where we replaced the Si tunnel barrier and capping layers with germanium while the triangular points are the samples that contain silicon. The black points are samples in which just a single layer of hyperdoped germanium is grown and the red points are the samples where we grew bilayer heterostructures. We observe a tendency for Si-containing samples to exhibit marginally increased carrier concentration and superconducting $T_{\textup{c}}$. A trend in $T_{\textup{c}}$ as a function of growth temperature is further demonstrated where we optimize the superconducting transition temperature at growth temperatures of 95-150C.
The scatter in $T_{\textup{c}}$ observed in Figure S3 for samples grown at similar conditions is attributed to small variations in atomic fluxes caused by drift during the course of film growth as well as the slight turbulence present within the Ge melt Lee et al. (1998). In principle, changing the Ge deposition source from an electron beam deposition source to a different technology would help mitigate these variations.
Figure S4: Hole concentration as a function of substrate temperature during growth. The same data point identification in Figure S3 is also used here. a Presents the entire set of samples, where we observe two outlier samples: one in which carrier concentration is drastically reduced compared to the other samples and one with significantly higher carrier concentration. We speculate poor dopant activation and Ga droplet formation as the cause of these two samples, respectively. We zoom in on the rest of the samples in b, where we see that a general downward trend is observed as substrate temperature increases. This trend is consistent with an increased growth temperature promoting Ga segregation, reducing the carrier concentration of the films.
Figure S5: Hall bar device fabrication and transport characterization of a bilayer heterostructure. a Presents an optical image of the fabricated device after wirebonding. We measure under current bias using a lock-in amplifier at low frequency to reduce noise in our measurements. b Shows the longitudinal resistance data upon cooldown of the device. These measurements are taken in a dilution refrigerator to ensure the sample remains superconducting down to mK temperatures with no re-entrant phase. However, we note that dilution fridges are non-ideal for measurements at higher temperatures where we see these vertical line artifacts during our measurement. This is caused by a delay in the thermometry of our fridge in updating the temperature of the mixing chamber plate. The transition temperature exists around 3 K, similar to the Van der Pauw measurement reported in Figure 1c on the same wafer. c Longitudinal resistance as a function of out-of-plane magnetic field also reports a critical magnetic field of the superconducting state to be $\sim$1 T, similar to what is reported for the Van der Pauw device presented in Figure S16a. d Hall trace on this device from which we calculate the carrier concentration to be $4.67\times 10^{21}$ cm-3, which is comparable to the Van der Pauw measurement of $4.15\times 10^{21}$ cm-3. From this data we conclude that the hyperdoped germanium thin film material is highly robust to typical photolithography and chlorine dry etch processing (more details are found in the Methods section). The inset schematic cartoons in b, c, and d represent the specific pins being measured for the presented traces based on the optical image shown in a. The dark boxes represent the pins being measured in these schematics.
S3 X-ray absorption measurement alignment and fitting
Figure S7: Comparison of the simulated fine structure derived from the CIF files of Ga:Ge interstitial system (DFT calculated system in Fig. 2a of main text) and amorphous Ga metal Sharma and Donohue (1962). Note that these Fourier transforms have been phase corrected, for simplicity. The arrows highlight the relatively large difference in experimental bond length of the nearest-neighbor scattering pair and those derived in the simulations.
S4 Impact of dopants on electronic structure
Figure S8: a Total density of states (DOS) calculated for the pure Ge crystal and the hyperdoped Ga:Ge shown in Fig. 2a of the main text, with Ga at interstitial (int.) and substitutional (sub.) lattice positions. The energies are aligned to the valence band maximum (VBM), set as the zero of the energy. The Fermi level in each case is indicated by a dashed vertical line. We observe further evidence that interstitial doping is unfavorable for for this system as it destroys the “clean” electronic bands of the host Ge semiconductor as evidenced by the disappearance of the parent Ge band gap and the emergence of metallic states. In contrast, substitutional Ga doping is expected to form acceptor states below the valence band edge, while the overall band structure near the gap remains largely unperturbed (see Figure 2d). b Total and projected DOS of the hyperdoped Ga:Ge crystal with substitutional fine-structure.
Figure S10: Fermi surface obtained after shifting the Fermi level by 40 meV below the value shown in Fig. 2d of the main text. a Contribution of bands i, ii, and iii to the 3D Fermi surface, with the color corresponding to the Fermi velocity. b (011) plane crossing the 3D Fermi surface. c 2D cut on this plane, showing the contribution of each band. The predicted narrow- or flat-band conditions are shown to be sensitive to the position of the Fermi level, as the pockets of high-mass carriers from bands $i$ and iii at the corners of the Brillouin Zone (R point) are shown to enlarge with a relatively minor shift in the Fermi level.
S5 Domain structure and near-surface Raman spectroscopy
Figure S11: a Low magnification image of the STEM lamella presented in Fig. 3 in the main text. We observe no domain structure as reported in previous reports of MBE-grown hyperdoped Ge films Strohbeen et al. (2023). The red dashed-line box is the region of higher magnification presented in b and Figure 3d.
Figure S12: a Raman Scattering penetration depth ($\mathrm{d_{\mathrm{RS}}}$) as a function of the laser excitation energy, determined by considering the absorption coefficients Aspnes and Studna (1983) for both the incoming ($\alpha_{\mathrm{out}}$) and the outgoing ($\alpha_{\mathrm{in}}$) light: $\mathrm{d_{\mathrm{RS}}}=1/(\alpha_{\mathrm{in}}+\alpha_{\mathrm{out}})\approx 1/\alpha_{in}$. b Comparison of Raman spectra recorded from Ga:Ge and Ge control MBE films at different excitation energies. At longer excitation wavelengths (785 nm), both the control and target films exhibit a comparable Raman response. Probing the surface-sensitive information with shorter wavelengths, the differences between the vibrational structure of the films become clearer. The appearance of disorder-activated LA modes below 200 cm-1 (from both Ge-type and Si-type sublattices) in the scattering volume yields comparable scattering intensities under these conditions for the control and doped sample. We suggest the origins of these bands are a combination of (local) lattice strain, phonon confinement and (partial) satisfaction of the resonance conditions of the Ge crystal $E_{1}+\Delta_{1}$ gap Alonso and Cardona (1988). Together, these features break the normal Raman scattering selection rules of a normal (001)-oriented bulk Ge film, i.e., resembling more the Raman spectra recorded using 785 nm laser light.
S6 Calculation of electron-phonon coupling
Figure S14: Eliashberg spectral function $\alpha^{2}F(\omega)$ calculated for substitutionally doped Ga:Ge and cumulative electron-phonon coupling strength $\lambda(\omega)$, showing that the coupling originates mostly from optical phonon modes. The total integrated $\lambda$ is 0.41, yielding a value of $T_{\textup{c}}=0.77$ K from the McMillan-Allen-Dynes formula [Eq. (2)].
To investigate the role of phonons as an underlying mechanism for the observed superconducting state, we perform additional calculations on the electron-phonon coupling strength within our model of hyperdoped Ge. Specifically, we compute the Eliashberg spectral function $\alpha^{2}F$ Giustino (2017), details of which are reported in the Methods section, to yield the total coupling strength, $\lambda$. From these quantities, the superconducting transition temperature, $T_{\textup{c}}$, is estimated using the McMillan-Allen-Dynes formula Allen and Dynes (1975); McMillan (1968). We calculate $\lambda=0.41$ and $T_{\textup{c}}=0.77$ K, providing evidence of a conventional phonon-mediated BCS-type pairing mechanism within this system. The Eliashberg spectral function reported in Figure S14 shows that this coupling originates primarily from the optical phonon modes, similar to what has been previously observed in the case of boron-doped diamond Ekimov et al. (2004) and silicon Bustarret et al. (2006).
S7 Synchrotron penetration depth and calculation of characteristic superconductor length scales
@@
Figure S17: Secondary-Ion Mass Specroscopy data for two samples investigated in this study, a PJS146 ($T_{\textup{c}}\sim$ 2.5 K) and b PJS147 ($T_{\textup{c}}\sim$ 3.5 K).
The magnetic field dependence of longitudinal resistance within the Josephson structure is presented in Figure S16a and b for out-of-plane and in-plane field configurations, respectively. At 300 mK we observe the out-of-plane $H_{\textup{c}}$ to be 0.79 T, decreasing to 0.14 T at 2 K. In-plane $H_{\textup{c}}$ is found to be 2.51 T, decreasing to 0.78 T at 2 K. We calculate the coherence length from this data using the following equation:
$\xi=\sqrt{\frac{\Phi_{\textup{0}}}{2\pi H_{\textup{c2}}(0)}}$ (3)
where $\Phi_{\textup{0}}$ is the magnetic flux quantum and $H_{\textup{c2}}(0)$ is the upper critical magnetic field extrapolated to zero Kelvin. London penetration depth is also calculated using:
$\lambda_{\textup{L}}=\sqrt{\frac{m^{*}}{2\pi n_{\textup{s}}q^{2}}}$ (4)
where $m^{*}$ is the effective mass of the hole carriers which we take to be the weighted average, $0.28m_{\textup{e}}$, and $n$ is the superfluid density which we obtain from our Hall measurements as $2n_{\textup{s}}=n_{\textup{h}}$. We calculate the penetration depth ($\lambda$) and coherence length ($\xi$) on the trilayer sample presented in Figures 3 and 4 to be 29 nm and 13.4 nm, respectively, similar to what has been previously observed in boron-doped diamond superconductors Bustarret et al. (2004).
S8 GIWAXS alignment, control, and stress analysis
[…]
@@
Figure S19: Integrated GIWAXS profile (qxyz) and structural refinement (Le Bail method) of the pure Ge epitaxial layer. Unstrained Ge crystal belongs to the high-symmetry cubic space group $Fd\overline{3}m$ (space group No. 227), characterized by equivalent lattice parameters ($a=b=c$) and a structural refinement of our undoped Si-Ge film aligns well with the reported bulk, cubic value ($a=5.657906\pm 0.0000009$ Å, Vol. = 181.1203 Å3Baker and Hart (1975).
Stress Analysis
The lattice mismatch is accommodated by an elastic strain in the hyperdoped epilayer, giving a biaxial stress of:
$\sigma=2\mu f\cdot(1+\nu)/(1-\nu)$ (5)
where $\mu$ is the shear modulus, $\nu$ Poisson’s ratio and $f$ the misfit parameter. Assuming the hyperdoped Ga:Ge epilayer forms a cubic unit cell when unconstrained, elastic isotropy is assumed in the relevant deformation directions and the misfit parameter is given by:
$f=\frac{(a_{\textup{Ga:Ge}}-a_{\textup{Ge}})}{a_{\textup{Ge}}}$ (6)
where $a_{\textup{Ga:Ge}}$ and $a_{\textup{Ge}}$ are the lattice parameters of the relaxed epilayer and substrate, respectively. We do not have such experimental data for the relaxed hyperdoped unit cell, however, since the epilayers studied here are pseudomorphically grown (Figure 3e), the values obtained from the structural refinement can be related to the relaxed lattice parameter of the Ga:Ge layer through:
$a_{\textup{Ga:Ge}}=a_{\textup{Ge}}[1+P(\Delta d/d)_{\perp}]$ (7)
where $P$ is the elastic parameter of the films McSkimin and Andreatch Jr. (1963). Regardless of the growth direction, the in-plane and out-of-plane biaxial strain components are given by:
$\epsilon_{\parallel}=(a_{\textup{Ga:Ge}}-a_{\textup{Ge}})/a_{\textup{Ge}}$ (8)
$\epsilon_{\perp}=\epsilon_{\parallel}(1-1/P)$ (9)
respectively. For our thin films grown on Ge (100); $P_{100}=C_{11}/(C_{11}+2C_{12})=0.5709$. To implement our calculation, we approximate the elastic constants $C_{ij}$ of the hyperdoped film to be comparable to the pure Ge crystal McSkimin and Andreatch Jr. (1963) (as there are no relevant data for GaGe alloys). Based on our hole carrier concentrations, the Ga concentration for the sample studied in Figure 3e is $\sim$11.22% and has a calculated relaxed (cubic) lattice parameter of $a_{\textup{Ga:Ge}}$ = 5.6614 Å, which is 0.11% larger than the Ge crystal.
S9 DFT and EXAFS data tables
@@
Table S2: EXAFS fitting parameters. Values are extracted from fits made to spectra recorded at 295 K using a Ga substitutional model. The lower ranked paths (in italics) are omitted from the actual fit model to avoid overfitting. The errors provided in the parentheses align with the last significant figures of the variable fit values. Further fitting details can be found in the Methods section. The path type can be tracked as: $S$ - single scattering pair; DS - double scattering; R - rattle.
[…]
@@
Data and materials availability:
Raw data available on Zenodo: doi.org/10.5281/zenodo.17065133 Strohbeen (2025)